# Retrieval-Enhanced Machine Learning: Synthesis and Opportunities

### v1

@@ 6.3.2. Decoupled Storage
Decoupled storage is defined as a storage method in REML systems where $g_{\omega}$ operates independent to $f_{\theta}$. In contrast to the coupled storage, where the entries are being updated dynamically online usually influenced by joint optimization of $f_{\theta}$ and $g_{\omega}$, decoupled storage involves offline population of entries, i.e., storage becomes read-only during the online stage. There are a few notable characteristics about REML systems with decoupled memory, including but not limited to:
@@
Since the retriever is completely independent from the predictive model, the implementation of the REML system becomes a lot easier compared to that with coupled storage (Borgeaud et al., 2022; Hui et al., 2022b; Yogatama et al., 2021; Khandelwal et al., 2020; Alon et al., 2022). When training the REML system, in decoupled storage architecture, one can either train the retriever separately or use an off-the-shelf retriever already publicly available (Shi et al., 2024). This also means that one can easily edit a REML system by simply replacing its components. Therefore, this design can guarantee a liberation from storage staleness and cold start problem unlike the coupled storage architecture. The ease of implementation stands out when the systems need to incorporate multiple storage that are multi-modal or multi-source (Fan et al., 2021; Yang et al., 2024a; Yu et al., 2022b), where it can be tricky to be implemented with coupled memory architecture.
@@ 7. Optimization
Optimization refers to how different model parameters are adjusted for performance. how does $g_{\omega}$ use $s\in\mathcal{S}$ provided by $f_{\theta}$ to update $\omega$? how does $f_{\theta}$ use $\mathcal{L}$ to update $\theta$?
@@ 7.1.1. No REML Optimization
Pre-trained language models (e.g., encoder representations) can be also used to produce latent representations for queries and documents, where simple similarity functions, such as dot product or cosine similarity, are used for computing relevance scores for a query-document pair. Even though these language models went through expensive optimization procedures, their optimizations are not REML-specific. Note that employing language model representations for retrieval with no optimization often do not perform well. For instance, Lien et al. (2024) demonstrated that using plain BERT or RoBERTa representations for zero-shot retrieval is substantially worse than term matching models, like BM25.
@@ 7.1.2. Independent Optimization
REML may take advantage of retrieval models whose optimization is independent of the predictive model’s parameters. Retrieval models are often trained using a set of query-document-relevance triplets. The relevance signal may come from (1) explicit annotations, e.g., from expert assessors or crowdworkers, (2) implicit feedback (Joachims, 2002; Joachims et al., 2017), e.g., user clicks, dwell time, mouse movements, etc., or (3) automatically generated weak signals (also known as distant supervision signals), such as appearance of a phrase, e.g., answer in the context of QA, in the documents, retrieval scores from another retrieval model (Dehghani et al., 2017), annotations produced by large language models (Thomas et al., 2023). Using these training triplets, retrieval models can be optimized using three different approaches: (1) pointwise ranking, (2) pairwise ranking, and (3) listwise ranking. Refer to Liu (2009) for more information on various ranking loss functions.
@@ 7.1.3. Conditional Optimization
(15) $\omega^{(t+1)}=\arg\min_{\theta}\frac{1}{|T|}\sum_{(x,y)\in T}\mathcal{L}\left(f_{\theta^{(t)}}\left(x;g_{\omega}\right),y\right)$
@@ 7.3.2. End-to-End Optimization
where both parameters sets $\theta$ and $\omega$ get updated simultaneously by optimizing an appropriate loss function for the downstream machine learning task. End-to-end optimization of REML, however, can be challenging. It is mostly due to the top $k$ item selection process of information access models in REML that makes the end-to-end REML model non-differentiable. Existing work make some simplifying assumptions to turn the optimization to a differentiable process. For instance, the RAG model from Lewis et al. (2020b) by marginalizing the retrieved document set to a set of pre-selected documents. A similar approach was later utilized by Sachan et al. (2021) for open-domain question answering. In addition to marginalization, RetGen (Zhang et al., 2022a) and EMDR2 (Singh et al., 2021) make a document independence assumption and computes the loss function as a summation over each individual document.
@@ 8. Evaluation
Evaluation refers to how retrieval components are benchmarked. evaluation constructs: extrinsic evaluation on the downstream task, intrinsic evaluation, attribution/support/explainability, efficiency given two retrieval models $g_{\omega}$ and $g_{\omega^{\prime}}$, how can we determine if $g_{\omega}\succ g_{\omega^{\prime}}$ intrinsically (e.g., using relevance judgments on retrieval results)? given two retrieval models $g_{\omega}$ and $g_{\omega^{\prime}}$, how can we determine if $g_{\omega}\succ g_{\omega^{\prime}}$ extrinsically (e.g., using L or $s\in\mathcal{S}$)?
@@ 8.1. Extrinsic evaluation
When evaluating a system extrinsically, we can pose hypotheses about relative system performance in several ways (Guu et al., 2020; Petroni et al., 2021; Lewis et al., 2020b). In non-overlapping system comparison, given two model tuples $\mathcal{M}=\langle f_{\theta},g_{\omega}\rangle$ and $\mathcal{M}^{\prime}=\langle f_{\theta^{\prime}},g_{\omega^{\prime}}\rangle$, determine if $\mathop{\mathbb{E}}[\mu(\mathcal{M}(x))]>\mathop{\mathbb{E}}[\mu(\mathcal{M}^{\prime}(x))]$. In fixed retrieval model comparison, given two model tuples $\mathcal{M}=\langle f_{\theta},g_{\omega}\rangle$ and $\mathcal{M}^{\prime}=\langle f_{\theta^{\prime}},g_{\omega}\rangle$, determine if $\mathop{\mathbb{E}}[\mu(\mathcal{M}(x))]>\mathop{\mathbb{E}}[\mu(\mathcal{M}^{\prime}(x))]$. As a special case, we can consider $g_{\omega^{*}}$ an optimal ranker according to some intrinsic criteria; this allows us to examine whether a system can effectively incorporate relevant items. In fixed predictive model comparison, given two model tuples $\mathcal{M}=\langle f_{\theta},g_{\omega}\rangle$ and $\mathcal{M}^{\prime}=\langle f_{\theta},g_{\omega^{\prime}}\rangle$, determine if $\mathop{\mathbb{E}}[\mu(\mathcal{M}(x))]>\mathop{\mathbb{E}}[\mu(\mathcal{M}^{\prime}(x))]$. In this case, we can consider $f_{\theta^{*}}$ as an optimal predictive model according to some intrinsic criteria; this allows us to examine whether a system can effectively retrieve relevant items.
@@ 8.2.2. Intrinsic evaluation of consumption
8.3. Datasets
Every REML system is tailored to execute a specific array of tasks. In literature, various benchmarks and datasets serve to assess these systems from diverse angles. Broadly, datasets fall into two categories: 1) those exclusively considering extrinsic evaluation of REML systems, assessing them solely based on end-to-end performance. 2) those furnishing retrieval relevance labels to evaluate retrieval performance in addition to end-to-end performance. That said, Table 5 illustrates the most commonly employed datasets and benchmarks in the literature.
@@ 9. Future Directions
Considering the extensive efforts made thus far in designing REML systems, this section presents new ideas for future work aimed at enhancing various aspects of these systems. Specifically, we propose several future directions for each of the previously discussed sections.
@@ 9.1.2. Retrieval System Aware Query Generation
Most of the existing query generation transformation and decomposition functions operate without taking into account the type and configuration of the retrieval model. However, using the query generation component that consider the specifications of the retrieval model can lead to the generation of more effective queries tailored to meet the model’s specific requirements. For instance, when using BM25 as the retriever model, which emphasizes term matching, queries with terms aligning closely with relevant documents can enhance retrieval performance. Conversely, in dense retrieval models that prioritize semantic similarity between queries and documents, generating queries with increased semantic alignment to relevant documents will be more advantageous.
@@ 9.2. Searching
9.2.1. Predictive Model Aware Retrieval Systems
Personalization literature in IR can be a good motivation for making a retrieval model that are tailored to a specific predictive model. For example, there can be a situation where multiple predictive models are being served by a few retrieval models (Salemi and Zamani, 2024b). In this case, the retrieval model can store meta-data of the predictive models they are serving, opening up opportunities to tailor the retrieval results to each or group of predictive models.
@@ 9.2.2. Redefining Relevancy
Related to the predictive model aware IR system, current IR models are built based on the relevance of a document. This relevancy is defined by whether this document is helpful in satisfying the users’ information need. However, in REML, the consumer of the documents are predictive models not human users (Salemi and Zamani, 2024a). Therefore, the relevancy should be redefined to incorporate the usefulness of the documents in predicting a targeted output.
@@ 9.3.2. Proactive REML
In practice, it is beneficial to not only answer the immediate question posed by the query, but also address potential follow up questions (Samarinas and Zamani, 2024; Liao et al., 2023). Follow up questions can transition to a new topic (e.g. purchasing a hotel after booking a flight), or dive deep on part of the initial answer (e.g. not only did Dave Grohl form the Foo Fighters, he was previously a drummer for Nirvana).
@@ 9.4. Storing
9.4.1. Shared collection
Figure 2 depicts a single $f_{\theta}$ interacting with multiple information access system. However, it is also possible to map multiple $f_{\theta_{i}}$ to a single information access system. In this specific situation, where multiple predictive model is sharing a single collection, an ability for models to learn what is and what is not important to store becomes critical as pushing irrelevant content to the shared storage can cause degradation in its usefulness and cause performance degradation of other $f_{\theta_{i}}$ sharing the same collection.
@@ 9.4.2. Storage staleness
9.5. Evaluation
9.5.1. Transparent Intrinsic evaluations
Although there have been a few attempts on systematic evaluation on REML models (Saad-Falcon et al., 2024; Es et al., 2024; Salemi and Zamani, 2024a), most of them use LLMs to perform evaluation. However, these metrics lack transparency in that they do not give valid reasoning why they are evaluated such a way. That said, researching transparent evaluation of REML systems is an important future direction.
9.5.2. Fairness, Trustworthiness, and Transparency of Output
Although these topics are rising, they are mainly evaluated on non-REML settings (e.g., text generation without retrieval-augmentation). Those evaluation criteria can pose different research direction tailored to REML applications. Identifying where the unfairness can happen (during retrieval and/or consumption) and checking if the unfairness propagates to the final output of the system can be a future research direction. Similar directions can be explored in model trustworthiness and transparency research.
@@ 9.6.4. One Information Access and Multiple Predictive Models
Most research focuses on developing a REML system for one task. On the other hand, information access systems can serve multiple predictive models, similar to the current search engines that serve billions of users, as studied in Salemi and Zamani (2024b). Optimizing information access components that provide service to multiple predictive models, aggregating and calibrating feedback across predictive models, and “personalizing” the retrieval result lists for each predictive model are important directions for future research.
@@ 10. Conclusion
In this work, we surveyed the current literature on retrieval-enhanced machine learning (REML) and synthesized it into consistent mathematical concepts, providing researchers with a formalized framework for REML. Additionally, by bridging information retrieval research and REML, we identify new opportunities and open avenues for future studies in this emerging research paradigm.

### v2

@@ 6.3.2. Decoupled Storage
Decoupled storage is defined as a storage method in REML systems where $g_{\omega}$ operates independent to $f_{\theta}$. In contrast to the coupled storage, where the entries are being updated dynamically online usually influenced by joint optimization of $f_{\theta}$ and $g_{\omega}$, decoupled storage involves offline population of entries, i.e., storage becomes read-only during the online stage. There are a few notable characteristics about REML systems with decoupled memory, including but not limited to the following.
@@
Since the retriever is completely independent from the predictive model, the implementation of the REML system becomes a lot easier compared to that with coupled storage (Borgeaud et al., 2022; Hui et al., 2022; Yogatama et al., 2021; Khandelwal et al., 2020; Alon et al., 2022). When training the REML system, in decoupled storage architecture, one can either train the retriever separately or use an off-the-shelf retriever already publicly available (Shi et al., 2024). This also means that one can easily edit a REML system by simply replacing its components. Therefore, this design can guarantee a liberation from storage staleness and cold start problem unlike the coupled storage architecture. The ease of implementation stands out when the systems need to incorporate multiple storage that are multi-modal or multi-source (Fan et al., 2021; Yang et al., 2024a; Yu et al., 2022b), where it can be tricky to be implemented with coupled memory architecture.
@@ 7.1.1. No REML Optimization
Pre-trained language models can be also used to produce latent representations for queries and documents, where simple similarity functions, such as dot product or cosine similarity, are used for computing relevance scores for a query-document pair. Even though these language models went through expensive optimization procedures, their optimizations are not REML-specific. Note that employing language model representations for retrieval with no optimization often do not perform well. For instance, Lien et al. (2024) demonstrated that using plain BERT or RoBERTa representations for zero-shot retrieval is substantially worse than term matching models, like BM25.
@@ 7.1.2. Independent Optimization
REML may take advantage of retrieval models whose optimization is independent of the predictive model’s parameters. Retrieval models are often trained using a set of query-document-relevance triplets. The relevance signal may come from (1) explicit annotations, e.g., from expert assessors or crowdworkers, (2) implicit feedback (Joachims, 2002; Joachims et al., 2017), e.g., user clicks, dwell time, mouse movements, etc, or (3) automatically generated weak signals (also known as distant supervision signals), such as appearance of a phrase, e.g., answer in the context of QA, in the documents, retrieval scores from another retrieval model (Dehghani et al., 2017), or annotations produced by large language models (Thomas et al., 2023). Using these training triplets, retrieval models can be optimized using three different approaches: (1) pointwise, (2) pairwise, or (3) listwise ranking. Refer to Liu (2009) for more information on various ranking loss functions.
@@ 7.1.3. Conditional Optimization
(15) $\omega^{(t+1)}=\arg\min_{\omega}\frac{1}{|T|}\sum_{(x,y)\in T}\mathcal{L}\left(f_{\theta^{(t)}}\left(x;g_{\omega}\right),y\right)$
@@ 7.3.2. End-to-End Optimization
where both parameters sets $\theta$ and $\omega$ get updated simultaneously by optimizing an appropriate loss function for the downstream machine learning task. End-to-end optimization of REML, however, can be challenging. It is mostly due to the top $k$ item selection process of information access models in REML that makes the end-to-end REML model non-differentiable. Existing work make some simplifying assumptions to turn the optimization to a differentiable process. For instance, the RAG model from Lewis et al. (2020) by marginalizing the retrieved document set to a set of pre-selected documents. A similar approach was later utilized by Sachan et al. (2021) for open-domain question answering. In addition to marginalization, RetGen (Zhang et al., 2022a) and EMDR2 (Singh et al., 2021) make a document independence assumption and computes the loss function as a summation over each individual document.
@@ 8.1. Extrinsic evaluation
When evaluating a system extrinsically, we can pose hypotheses about relative system performance in several ways (Guu et al., 2020; Petroni et al., 2021; Lewis et al., 2020). In non-overlapping system comparison, given two model tuples $\mathcal{M}=\langle f_{\theta},g_{\omega}\rangle$ and $\mathcal{M}^{\prime}=\langle f_{\theta^{\prime}},g_{\omega^{\prime}}\rangle$, determine if $\mathop{\mathbb{E}}[\mu(\mathcal{M}(x))]>\mathop{\mathbb{E}}[\mu(\mathcal{M}^{\prime}(x))]$. In fixed retrieval model comparison, given two model tuples $\mathcal{M}=\langle f_{\theta},g_{\omega}\rangle$ and $\mathcal{M}^{\prime}=\langle f_{\theta^{\prime}},g_{\omega}\rangle$, determine if $\mathop{\mathbb{E}}[\mu(\mathcal{M}(x))]>\mathop{\mathbb{E}}[\mu(\mathcal{M}^{\prime}(x))]$. As a special case, we can consider $g_{\omega^{*}}$ an optimal ranker according to some intrinsic criteria; this allows us to examine whether a system can effectively incorporate relevant items. In fixed predictive model comparison, given two model tuples $\mathcal{M}=\langle f_{\theta},g_{\omega}\rangle$ and $\mathcal{M}^{\prime}=\langle f_{\theta},g_{\omega^{\prime}}\rangle$, determine if $\mathop{\mathbb{E}}[\mu(\mathcal{M}(x))]>\mathop{\mathbb{E}}[\mu(\mathcal{M}^{\prime}(x))]$. In this case, we can consider $f_{\theta^{*}}$ as an optimal predictive model according to some intrinsic criteria; this allows us to examine whether a system can effectively retrieve relevant items.
@@ 8.2.2. Intrinsic evaluation of consumption
8.3. Datasets and Benchmarks
REML systems can be evaluated on various tasks. In literature, various benchmarks and datasets serve to assess these systems from diverse angles. Broadly, datasets fall into two categories: 1) those exclusively considering extrinsic evaluation of REML systems, assessing them solely based on end-to-end performance. 2) those furnishing retrieval relevance labels for intrinsic evaluation in addition to end-to-end performance. Table 5 illustrates the most commonly employed datasets and benchmarks in the literature.
@@ 9. Future Directions
To enhance REML systems, we propose future work for each of the previously discussed sections.
@@ 9.1.2. Retrieval System Aware Query Generation
Most query generation and decomposition functions overlook the type and configuration of the retrieval model, but aligning queries with the model’s specifics can enhance performance. For instance, BM25, which emphasizes exact term matching, performs better with queries that closely align with terms in relevant documents. On the other hand, dense retrieval models, which focus on semantic similarity, benefit more from queries that are semantically aligned with the content of relevant documents. Tailoring query generation to the retrieval model ensures the queries meet the model’s unique requirements, improving retrieval effectiveness.
@@ 9.2. Searching
9.2.1. Predictive Model-Aware Retrieval Systems
Approaches like search personalization could be used for tailoring retrieval results to a specific predictive model. For example, there can be a situation where multiple predictive models are being served by a few retrieval models (Salemi and Zamani, 2024b). In this case, the retrieval model can store meta-data of the predictive models they are serving, opening up opportunities to tailor the retrieval results to each predictive model.
@@ 9.2.2. Redefining Relevancy
In predictive model-aware retrieval systems, IR models usually evaluate a document’s relevance based on how well it meets a human user’s information need. However, in REML, the primary consumer is the predictive model, not the human user. Research shows that utility judgments—assessing a document’s usefulness for a specific task—correlate more strongly with the REML model’s performance than traditional relevance labels (Salemi and Zamani, 2024a; Kim and Diaz, 2024; Zhang et al., 2024b). This underscores the need for further research on how to best represent relevance in the context of REML.
@@ 9.3.2. Proactive REML
In practice, it is beneficial to not only answer the immediate question posed by the query, but also address potential follow up questions (Samarinas and Zamani, 2024; Liao et al., 2023). Follow up questions can transition to a new topic (e.g. purchasing a hotel after booking a flight), or dive deep on part of the initial answer.
@@ 9.4. Storing
9.4.1. Shared Storage
Figure 2 depicts a single $f_{\theta}$ interacting with multiple information access system. However, it is also possible to map multiple $f_{\theta_{i}}$ to a single information access system. In this scenario, where multiple predictive model is sharing a single collection, the ability of the models to learn what is important to store becomes crucial as pushing irrelevant content to the shared storage can cause degradation in its usefulness and cause performance degradation of other $f_{\theta_{i}}$ sharing the same collection. This direction is particularly valuable to explore in multi-agent settings (Park et al., 2023).
@@ 9.5.4. One Information Access and Multiple Predictive Models
Most prior work focuses on developing a REML system for one task. On the other hand, information access systems can serve multiple predictive models, similar to the current search engines that serve billions of users (Salemi and Zamani, 2024b). Optimizing information access components that provide service to multiple predictive models, aggregating and calibrating feedback across predictive models, and “personalizing” the retrieval result lists for each predictive model are important future directions.
9.6. Evaluation
9.6.1. Broadening Evaluation Approaches
Evaluations of REML systems predominantly focus on extrinsic measures of system effectiveness, and studies that conduct both extrinsic and intrinsic evaluations mainly employ reference-based methods with relevance labels. Researchers could broaden this scope by assessing additional qualities of REML systems, such as citation quality (Gao et al., 2023b), robustness (Yoran et al., 2024), and context utilization (Ru et al., 2024). Additionally, an often overlooked aspect is the evaluation of trustworthiness and societal impact; although recognized as important, these factors are typically discussed in non-REML contexts like pure information retrieval or text generation. In REML applications, trustworthiness evaluation might present new research directions. For instance, existing definitions of fairness may require revision within the REML context (Kim and Diaz, 2024), and the propagation of related issues through retrieval and consumption processes may need re-examination. Regarding transparency in evaluation, most model-based, reference-free automatic evaluations of REML models rely on large language models (Saad-Falcon et al., 2024; Es et al., 2024; Salemi and Zamani, 2024a; Guinet et al., 2024), which often lack transparency because they do not provide clear reasoning for their assessments. These automatic evaluation models can be meta-evaluated based on measurement theory (Hand, 2010) for further validation.
@@ 10. Conclusion
In this work, we survey the current literature on retrieval-enhanced machine learning (REML) and synthesize it into consistent mathematical concepts, providing researchers with a formalized framework for REML. By bridging the gap between information retrieval research and REML, we also identify new opportunities and open avenues for future studies in this emerging research paradigm.